OpenStudy (anonymous):

messaging plan A 35.00 monthly and fee per message is 0.10 . Plan B say that there monthly fee is 18.00 and there fee per message is 0.30 for what number of text messages will both plans cost the same

OpenStudy (anonymous):

a.55 b.85 c.110 d.140

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

You just have to plug it in. .30 times what is 55 and .10 times what is 55? If they match, then that is your answer. Still need help? :o)

OpenStudy (anonymous):

No wait, thats wrong. It should be .30*55= ?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

Okay, well for .30, 55 text messages will be $16.50, for .10, 55 text messages will be $5.50. For .30, 85 text messages will be $25.50. For .10, 85 text messages will be $8.50, etc.

OpenStudy (anonymous):

I'm not sure how to solve it.. sorry.

OpenStudy (anonymous):

ok thanks for trying

OpenStudy (anonymous):

Hi! Let us say the number of text messages (on each plan) that would cost same for both plans be x. Therefore, cost due to the first plan would be: 35(monthly payment irrespective of number of texts) + 0.10 * x (for the number of texts) = 35+ 0.10*x Now cost due to second plan would be =18 + 0.30*x It is given that these two costs are equal 35+0.10x = 18 + 0.30x Therefore 0.20x = 35-18 = 17 (2/10) x = 17; Therefore x = 17*10/2 = 17*5 = 85 Therefore, if u send 85 messages on each plan, both plans would cost the same